# Bayesian WeakS-to-Strong from Text Classification to Generation

### v1

@@ 1 Introduction
With the increase in computing power and the amount of training data available, the capabilities of large language models (LLMs) have been continuously brought closer to humans in many aspects. Despite their impressive performance, the preferences and values of pre-trained LLMs do not always align with humans, and dedicated approaches are needed to tackle the problem. Based on large-scale instruction datasets, supervised finetuning (SFT) encourages LLMs to follow human instructions more strictly and respond more safely [28]. Reinforcement learning (RL) is commonly applied to such alignment. By collecting model output values and the corresponding human feedback, the model can be finetuned by RL to avoid generating undesirable outputs [33, 3, 18, 17, 1].
@@
Since no current model has surpassed human intelligence yet, existing alignment methods, such as SFT and RL from human feedback (RLHF), remain effective. However, it is worthwhile considering future scenarios where artificial intelligence (AI) might surpass human intelligence in all aspects. Would the current alignment methods still be effective for such super AI models? How could humans supervise the super AI? To simulate this future scenario, an analogy situation is designed that downgrades both sides: using a weak model to simulate humans and a strong model to simulate future super AI [7], which is termed as superalignment. It has been demonstrated that adding a simple auxiliary loss can achieve effective Weak-to-Strong generalization, even if the weak model’s supervision contains many errors, which offers hope of achieving superalignment. Nonetheless, this is just the beginning of exploring along the path of Weak-to-Strong.
@@
This paper extends the discussion on Weak-to-Strong in two directions. First, given the inherent capability gap between the weak model and the strong model, we propose using an ensemble of multiple weak models to improve the quality of weak supervision, which is called WeakS-to-Strong. This also accounts for the scenario where human opinions might diverge in tasks without a commonly accepted standard. Several approaches have been studied to effectively leverage the diversity of different weak models, and we adapt a Bayesian approach referred to as evidential deep learning (EDL) [24] to better estimate broader human preferences by learning a prior distribution over the weak labels produced by the weak models. Furthermore, the Weak-to-Strong task was primarily studied for text classification tasks [7]. This paper extends the scope to text generation tasks and demonstrates that the proposed Bayesian WeakS-to-Strong approach is effective for both. To better align with human preferences, a variant of direct preference optimization (DPO) [22] called conservative DPO (cDPO) [9] is used to finetune the strong model further on RL principles.
@@
When applied to text generation, a token-level confidence measure is proposed to achieve soft labels for strong model training. We also propose the modified DPO algorithm under the Bayesian WeakS-to-Strong framework to further improve text generation performance.
@@ 2 Related Work
AI Alignment. Aligning LLMs with human preferences has been a long-standing goal. Instruction tuning uses extensive datasets to improve LLMs’ adherence to human instructions [28]. RL allows LLMs to learn what types of responses humans prefer or dislike, with proximal policy optimization (PPO) being an effective RL method first applied to LLMs and becoming part of the standard RLHF process [33, 3, 18, 17, 1]. However, PPO training can be unstable, leading to the development of DPO [22]. Due to the high cost of obtaining human preference data, there is now exploration into using LLMs to simulate human preferences, provide feedback, and finetune models [13, 4, 10].
Variability in Human Opinions. In the process of aligning AI with human preferences, it is important to consider the inconsistency of human preferences [14], which often leads to multi-label problems. Previously, many approaches used simple methods like voting, aggregation, and averaging to handle multi-labels [8, 16, 19, 20]. However, these methods do not effectively capture the preference differences of individual annotators included in the multiple labels. To better estimate the diversity of human preferences, Bayesian principles have been introduced. Deep neural network models can be used to predict prior distributions, which are considered to produce the multiple available labels to estimate a broader range of human preferences [24, 31, 30].
Weak-to-Strong. The goal of Weak-to-Strong is to use a weak model to better supervise a strong model. OpenAI demonstrated that adding auxiliary confidence loss from the strong model itself can significantly improve the Weak-to-Strong performance [7]. Following OpenAI’s work, several studies emerged to introduce multiple weak models, used either in series or parallel, to improve the quality of supervision provided by the weak models [15, 23]. In addition, confidence scores are incorporated to help the strong model assess the supervision quality provided by the weak models [11] and weak model can be directly used to modify the output of the strong model [12].
@@ 3.1 Preliminary: Weak-to-Strong
The Weak-to-Strong pipeline [7] involves three steps: (i) create a weak supervisor by finetuning a small pre-trained model on ground-truth labels; (ii) train a strong student model with weak supervision by finetuning a pre-trained LLM using “weak labels” generated by weak supervisors, where is the parameters of the strong model; (iii) finetune the large pre-trained model directly using ground-truth labels which serves as the ceiling.
To leverage the superior generalization capabilities and prior knowledge of strong model, a loss function with auxiliary confidence loss is proposed [7]:
where represents the weak label from weak model, refers to the predicted class of strong model given input , and denotes the cross-entropy loss. The second term is an (optional) auxiliary self-training loss designed to increase the confidence of the strong model to itself. The weight of the second loss, , linearly grows up from 0 to a pre-defined hyper-parameter , which gradually reduces the weight on the weak labels and increase the weight on self-training when the number of training step increases.
@@ 3.2 Extending Weak-to-Strong with multiple weak models
Although it has been shown that the Weak-to-Strong approach can recover part of the strong model’s performance [7], the errors in weak labels limit the performance of Weak-to-Strong generalization. In response to this problem, we propose to leverage the complementarity of the error patterns of multiple weak models using the ensemble strategy, which is referred to as WeakS-to-Strong.
@@
where is the number of weak models, is the th weak label produced by the th weak model, and is a pre-defined weight of the loss regarding the the th weak model. This approach is referred to as a naive multi-weak system in the rest of the paper.
@@ 3.3 Bayesian WeakS-to-Strong
For superalignment, multiple weak models are used to mimic the subjective preferences of multiple humans, which can be considered as observations drawn from an underlying distribution of the opinions of all humans. The naive approach described in Section 3.2 solely relies on these observations. The number of observations (human annotations or weak labels) is often very limited due to the considerable cost of hiring a new human annotator or training a new weak model. Such a limited number of observations may not result in a good approximation of the true human opinion distribution. Having biased preferences or values is particularly unacceptable in the safety domain and can cause a failure of superalignment. Therefore, we propose a Bayesian WeakS-to-Strong approach based on EDL [24] to estimate the human opinion distribution based on the weak labels. Figure 1(c) illustrates the framework where three weak models are involved.
Consider a weak label from the weak model , which is a one-hot vector with being one if it belongs to class and zero otherwise. is sampled from a categorical distribution (of the labels of weak models) , where each component corresponds to the probability assignment over the possible classes EDL places a Dirichlet prior over the categorical distribution representing the probability of each categorical probability assignment, hence modelling second-order probability , where is the hyperparameter of the Dirichlet prior. The strong model is trained to predict for each input by minimizing the negative log-likelihood of sampling given the predicted Dirichlet prior:
where is the number of classes, is the Dirichlet strength, and is the value of label . Following [24], a regularization term (see Appendix B for details) is added to calibrate uncertainty estimation, resulting in the EDL loss where is the coefficient.
@@ 4 WeakS-to-Strong for Sequence Generation
4.1 Label Formulation via Joint-Decoding
To enable the strong model to directly generate trustworthy content rather than only being trained to understand whether the content is trustworthy or not, we propose to extend the scope of Weak-to-Strong from text classification to text generation. We also propose to jointly decode with multiple weak models to derive the sequence-level training target of the strong model. In contrast to the Naive multi-weak scheme, joint decoding employs multiple weak models to collaboratively determine one single target, reducing the risk of the strong model being affected by the potential biases that exist in the label space of the separate weak models.
Specifically in this paper, we perform joint decoding in a re-ranking fashion. For each weak model, the top output sequences are generated by beam search in decoding. We gather the output sequences from the weak models to form a list of sequences. Next, for each sequence, scores are computed by generating it using each weak model separately (via teacher forcing) and aggregating the output (log-)probabilities. We perform a weighted sum of the scores as the final score assigned to each sequence, where the same set of weights is fixed within each task. The sequence with the highest final score will be used as the target sequence for strong model training. That is,
where is the th output sequence, is the weight assigned to the th weak model, represents model parameters of the th weak model, and .
4.2 Confidence measure for weak sequence labels
@@
The key challenge of directly applying the Weak-to-Strong loss [7] to the sequence generation task is the token-level soft labelling for the target sequence. As the tokenizers are different between the weak and the strong models, it is infeasible to obtain a one-to-one mapping from weak model output distributions to each token in the target sequence.
@@
To obtain the soft label for the strong model using weak model output probabilities and bridge the gap caused by different tokenizers, we use words as an intermediary. Figure 2 shows an example of the process in three stages. In stage 1, the per-token output probabilities of the weak models are obtained when generating output sequences. The probabilities of wordpieces in a word, referred to as wordpiece scores, are then multiplied together to obtain the score of word . In stage 2, the word score is used to assign probabilities to tokens from the strong model tokenizer, following Eqn. (7).
where we use a word containing two wordpieces from both weak and strong tokenizers as an example, and and are both token strings that can form , which are generated by the tokenizers of the weak and strong models respectively. To obtain the actual assignment of scores to each strong model token , instead of assigning equal probabilities to all tokens involved, we compute a weight from the strong model confidence when predicting and to approximate the decomposition by
@@
where is strong model confidence at the step predicting wordpiece , which is the maximum probability in the strong model output distribution at that step in practice. In this way, smaller target confidence scores are allocated to tokens with lower strong model confidence (i.e., harder to predict), while larger confidence scores are allocated to tokens with higher strong model confidence.
@@
After obtaining the probability of the target token of the strong model, the probabilities of other tokens can be obtained by scaling strong prediction logits, which can be treated as the soft label, as shown in Figure 2. Then the obtained soft labels can be handled using methods similar to those used in classification. Notably, during the computation of EDL loss, the sparsity caused by high dimensional spaces results in a large KL penalty term. To solve this problem, a coefficient is added to the KL penalty to balance it with the magnitude of the negative log-likelihood term. Additionally, clamping is applied to restrict all values within an appropriate range, preventing potentially extremely large outliers on any particular token.
@@ 4.3 DPO for Sequence Generation Optimization
Different from classification tasks, sequence generation tasks often benefit from sequence-level objectives that directly optimize the entire sequence jointly rather than the individual tokens separately. To further improve the strong model for sequence generation, direct preference optimization (DPO) [22] is investigated for WeakS-to-Strong after supervised finetuning, where we propose to use weak models to provide the preference for the strong model generation.
The standard DPO loss [22] can be written as follows:
where denotes the model to be optimized with parameter , refer to the reference model, initialized as supervised finetuned model, and , refer to preferred and dispreferred response pair respectively. Intuitively, the loss function increases the likelihood of the preferred completions and decreases the likelihood of dispreferred completions, scaled by , controlling the deviation from reference model .
In practice, a strong model pre-trained by supervised finetuning generates output sequences based on a given input, which are scored by the weak models in the same way as scoring in joint decoding is performed (see Section 4.1). The sequence with the highest score is viewed as the preferred sequence, and that with the lowest score is the dispreferred sequence, as shown below as
where is the th output sequence generated by the strong model, and is computed using the th weak model. The dispreferred sequence can be computed similarly by
Considering potential errors by weak models, a variant of DPO, conservative DPO (cDPO) [9] with a more conservative target distribution is applied in our work. The loss of cDPO is
where is a small constant probability that labels are flipped to make DPO more conservative.
@@ 5.1 Datasets and Models
Classification Task. The setup of the classification task follows [7]. The SciQ dataset [29] is used, which contains 13,679 crowdsourced science exam questions about Physics, Chemistry and Biology, among others. The questions are in multiple-choice format with 4 answer options each. In our experiment, 5k data samples were extracted for training weak models and another 5k samples were reserved for generating weak labels to train the strong model. The standard test set which contains 1k data samples was used for the test. The data is restructured into a balanced binary classification task, i.e., given a question and an answer, the model is required to determine whether the answer is correct.
For classification, the Qwen-7B [2]222https://huggingface.co/Qwen/Qwen-7B model was applied as the strong student model. Three models were used as weak teachers: GPT2-Large [21]333https://huggingface.co/openai-community/gpt2-large, OPT-1.3B [32]444https://huggingface.co/facebook/opt-1.3b and Pythia-1.4B [6]555https://huggingface.co/EleutherAI/pythia-1.4b. The last linear layer which maps the embeddings to tokens is replaced with a linear classification head with two outputs to adapt language models to the classification setting.
@@
Slot filling. The performance of WeakS-to-Strong on the reliability of generated content was evaluated on the slot filling task, which is a crucial spoken language understanding task aiming at filling in the correct value for predefined slots (e.g., restaurant and hotel names). SLURP dataset [5] was used which contains 16.5k sentences and 72k audio recordings of single-turn user interactions with a home assistant, annotated with scenarios, actions and entities. Only the reference transcriptions of the speech were used for training. Following [25] and [26], we designed the prompt with slot keys and descriptions in the same way. In our setup, 2k utterances from the train split were extracted for training the weak models, and another 2k utterances were reserved for generating weak labels and training the strong model. We report the performance of both weak and strong models on the standard SLURP test set.

### v2

@@ 1 Introduction
With the increase in computing power and the amount of training data available, the capabilities of large language models (LLMs) have been continuously brought closer to humans in many aspects. Despite their impressive performance, the preferences and values of pre-trained LLMs do not always align with humans, and dedicated approaches are needed to tackle the problem. Based on large-scale instruction datasets, supervised finetuning (SFT) encourages LLMs to follow human instructions more strictly and respond more safely (Wei et al., 2022). Reinforcement learning (RL) is commonly applied to such alignment. By collecting model output values and the corresponding human feedback, the model can be finetuned by RL to avoid generating undesirable outputs (Ziegler et al., 2019; Bai et al., 2022a; Ouyang et al., 2022; Nakano et al., 2021; Askell et al., 2021).
@@
Since no current model has yet surpassed human intelligence, alignment methods, such as SFT and RL from human feedback (RLHF), remain effective. However, it is worthwhile considering future scenarios where artificial intelligence (AI) might surpass human intelligence in all aspects. Would the current alignment methods still be effective for such super AI models? How could humans supervise the super AI? To simulate this future scenario, an analogy situation is designed that downgrades both sides: using a weak model to simulate humans and a strong model to simulate future super AI (Burns et al., 2023), which is termed as superalignment. It has been demonstrated that adding a simple auxiliary loss can achieve effective Weak-to-Strong generalization, even if the weak model’s supervision contains many errors, which offers hope of achieving superalignment. Nonetheless, this is just the beginning of exploring along the path of Weak-to-Strong.
@@
This paper extends the discussion on Weak-to-Strong in two directions. First, given the inherent capability gap between the weak model and the strong model, we propose using an ensemble of multiple weak models to improve the quality of weak supervision, which is called WeakS-to-Strong. This also accounts for the scenario where human opinions might diverge in tasks without a commonly accepted standard. Several approaches have been studied to effectively leverage the diversity of different weak models, and we adapt a Bayesian approach referred to as evidential deep learning (EDL) (Sensoy et al., 2018) to better estimate broader human preferences by learning a prior distribution over the weak labels produced by the weak models. Furthermore, the Weak-to-Strong task was primarily studied for text classification tasks (Burns et al., 2023). This paper extends the scope to text generation and compares three different approaches, Naive Multi-Weak, Joint Decoding and Bayesian Multi-Weak, as shown in Figure 1. The proposed Bayesian WeakS-to-Strong approach is demonstrated effective for both classification and generation. To better align with human preferences, a variant of direct preference optimization (DPO) (Rafailov et al., 2023) called conservative DPO (cDPO) (Eric, 2023) is used to finetune the strong model further on RL principles.
@@
When applied to text generation, a token-level probability estimation is proposed to achieve soft labels for strong model training. We also propose the modified DPO algorithm under the Bayesian WeakS-to-Strong framework to further improve text generation performance.
@@ 2 Related Work
AI Alignment. Aligning LLMs with human preferences has been a long-standing goal. Instruction tuning uses extensive datasets to improve LLMs’ adherence to human instructions (Wei et al., 2022). RL allows LLMs to learn what types of responses humans prefer or dislike, with proximal policy optimization (PPO) being an effective RL method first applied to LLMs and becoming part of the standard RLHF process (Ziegler et al., 2019; Bai et al., 2022a; Ouyang et al., 2022; Nakano et al., 2021; Askell et al., 2021). However, PPO training can be unstable, leading to the development of DPO (Rafailov et al., 2023). Given the high cost of obtaining human preference data, researchers are now exploring the use of LLMs to simulate human preferences, provide feedback, and finetune models (Lee et al., 2023; Bai et al., 2022b; Gulcehre et al., 2023).
Variability in Human Opinions. In the process of aligning AI with human preferences, it is important to consider the inconsistency of human preferences (Liu et al., 2023), which often leads to multi-label problems. Previously, many approaches used simple methods like voting, aggregation, and averaging to handle multi-labels (Davani et al., 2022; Munos et al., 2023; Paun & Simpson, 2021; Prabhakaran et al., 2021). However, these methods do not effectively capture the preference differences of individual annotators included in the multiple labels. To better estimate the diversity of human preferences, Bayesian principles have been introduced. Deep learning models can be used to predict prior distributions, which are considered to produce the multiple available labels to estimate a broader range of human preferences (Sensoy et al., 2018; Wu et al., 2022; 2023).
Weak-to-Strong. The goal of Weak-to-Strong is to use a weak model to better supervise a strong model. OpenAI demonstrated that adding auxiliary confidence loss from the strong model itself can significantly improve the Weak-to-Strong performance (Burns et al., 2023). Following OpenAI’s work, several studies emerged to introduce multiple weak models, used either in series or parallel, to improve the quality of supervision provided by the weak models (Liu & Alahi, 2024; Sang et al., 2024). Early model ensemble methods like Adaboost (Freund & Schapire, 1995) and Bootstrap aggregating (Leo, 1996) were explored in these works. Furthermore, confidence scores are incorporated to help the strong model assess the supervision quality provided by the weak models (Guo et al., 2024) and the weak model can be directly used to modify the output of the strong model (Ji et al., 2024).
@@ 3.1 Preliminary: Weak-to-Strong
The Weak-to-Strong pipeline (Burns et al., 2023) involves three steps: (i) create a weak supervisor by finetuning a small pre-trained model on ground-truth labels; (ii) train a strong student model with weak supervision by finetuning a pre-trained LLM using “weak labels” generated by weak supervisors, where is the parameters of the strong model; (iii) finetune the large pre-trained model directly using ground-truth labels which serve as the ceiling.
To leverage the superior generalization capabilities and prior knowledge of the strong model, a loss function with auxiliary confidence loss is proposed (Burns et al., 2023):
where represents the weak label from weak model, refers to the predicted class of strong model given input , and denotes the cross-entropy loss. The second term is an (optional) auxiliary self-training loss designed to increase the confidence of the strong model in itself. The weight of the second loss linearly grows up from 0 to a pre-defined hyper-parameter , which gradually reduces the weight on the weak labels and increases the weight on self-training when the number of training steps increases.
@@ 3.2 Extending Weak-to-Strong with multiple weak models
Although it has been shown that the Weak-to-Strong approach can recover part of the strong model’s performance (Burns et al., 2023), the errors in weak labels limit the performance of Weak-to-Strong generalization. In response to this problem, we propose to leverage the complementarity of the error patterns of multiple weak models using an ensemble strategy, which is referred to as WeakS-to-Strong.
@@
where is the number of weak models, is the th weak label produced by the th weak model, and is a pre-defined weight of the loss regarding the the th weak model. This approach is referred to as a Naive Multi-Weak system in the rest of the paper (as illustrated in Figure 1(a)), which is treated as one of the baselines.
@@ 3.3 Bayesian WeakS-to-Strong
For superalignment, multiple weak models are used to mimic the subjective preferences of multiple humans, which can be considered as observations drawn from an underlying distribution of the opinions of all humans. The naive approach described in Section 3.2 solely relies on these observations. The number of observations (human annotations or weak labels) is often very limited due to the considerable cost of hiring a new human annotator or training a new weak model. Such a limited number of observations may not result in a good approximation of the true human opinion distribution. Having biased preferences or values is particularly unacceptable in the safety domain and can cause a failure of superalignment. Therefore, we propose a Bayesian WeakS-to-Strong approach based on EDL (Sensoy et al., 2018) to estimate the human opinion distribution based on the weak labels. Figure 1(c) illustrates the framework where three weak models are involved.
Consider a weak label from the weak model , which is a one-hot vector with being one if it belongs to class and zero otherwise. is sampled from a categorical distribution of weak labels , where each component corresponds to the probability assignment over the possible classes EDL places a Dirichlet prior over the categorical distribution representing the probability of each categorical probability assignment, hence modelling second-order probability , where is the hyperparameter of the Dirichlet prior. The strong model is trained to predict for each input by minimizing the negative log-likelihood of sampling given the predicted Dirichlet prior:
where is the number of classes, is the Dirichlet strength, and is the value of label . Following Sensoy et al. (2018), a regularization term (see Appendix G for details) is added to calibrate uncertainty estimation, resulting in the EDL loss where is the coefficient.
@@ 4 WeakS-to-Strong for Sequence Generation
4.1 Probability estimation for weak sequence labels
@@
To enable the strong model to directly generate trustworthy content rather than only being trained to understand whether the content is trustworthy or not, we propose to extend the scope of Weak-to-Strong from text classification to text generation. The key challenge of directly applying the Weak-to-Strong loss to the sequence generation task is the token-level soft labelling for the target sequence. As the tokenizers are different between weak and strong models, it is infeasible to obtain a one-to-one mapping from weak model output distributions to each token in the target sequence.
To obtain the soft label for the strong model using weak model output probabilities and bridge the gap caused by different tokenizers, we use words as an intermediary, following the equation
@@
where we use a word containing two wordpieces from both weak and strong tokenizers as an example, and and are both token strings that can form word , which are generated by the tokenizers of the weak and strong models respectively. Figure 2 shows an example of the process in three stages. In stage 1, the per-token output probabilities of weak models are obtained when generating output sequences. The probabilities of wordpieces in a word are then multiplied together to obtain the score of word , following Eqn. (6).
In stage 2, the word probability is used to assign probabilities to tokens from the strong model tokenizer. In the process of training the strong model via teacher-forcing, the model gives a probability to each token by applying softmax to the output logits. This probability can be seen as the model’s confidence in predicting the target token, where we assume that the weak model and strong model have similar confidence for wordpiece tokens in similar positions. This allows us to to obtain the actual assignment of scores to each target token , instead of assigning equal probabilities to all tokens involved. Taking an example of splitting a word into two target wordpieces and , the decomposition can be approximated by
@@
where is strong model confidence at the step predicting wordpiece that is the maximum probability in the strong model output distribution at that step in practice. In this way, lower target probabilities are allocated to tokens with lower strong model confidence, while higher probabilities are allocated to tokens with higher strong model confidence.
@@
After obtaining the probability of the target token of the strong model, the probabilities of other categories can be obtained by scaling strong prediction probabilities, which can be treated as the soft label, as shown in Figure 2. Then the obtained soft labels can be handled using methods similar to those used in classification (as described in Section 3). Notably, during the computation of EDL loss, the sparsity caused by high dimensional spaces results in a large KL (Kullback–Leibler) penalty term. To solve this problem, a coefficient is added to the KL penalty to balance it with the magnitude of the negative log-likelihood term. Additionally, clamping is applied to restrict all values within an appropriate range, preventing potentially extremely large outliers on any particular token.
@@ 4.2 DPO for sequence generation optimization
Different from classification tasks, sequence generation tasks often benefit from sequence-level objectives that directly optimize the entire sequence jointly rather than the individual tokens separately. To further improve the strong model for sequence generation, direct preference optimization (DPO) (Rafailov et al., 2023) is investigated for WeakS-to-Strong after supervised finetuning, where we propose to use weak models to provide the preference for the strong model generation.
After the strong model is pretrained by supervised finetuning, it generates output sequences based on a given output. Then for each sequence, scores are computed by generating it using weak models separately (via teacher forcing) and aggregating the output (log-)probabilities. A weighted sum of the scores is performed as the final score assigned to each sequence. The sequence with the highest final score is viewed as the preferred sequence in DPO training, as shown below
where is the th output sequence generated by the strong model, is computed using the th weak model with model parameters , and is the weight assigned to the th weak model. The dispreferred sequence can be computed similarly by
Considering potential errors by weak models, a variant of DPO, conservative DPO (cDPO) (Eric, 2023) with a more conservative target distribution is applied in our work. The loss of cDPO is
where is a small constant probability that labels are flipped to make DPO more conservative, and is the standard DPO loss in (Rafailov et al., 2023), which can be written as
@@ 5.1 Datasets
Classification Task. The setup of the classification task follows Burns et al. (2023). The SciQ dataset (Welbl et al., 2017) is used, which contains 13,679 crowdsourced science exam questions about Physics, Chemistry and Biology, among others. The questions are in multiple-choice format with 4 answer options each. In our experiment, 5k data samples were extracted for training weak models and another 5k samples were reserved for generating weak labels to train the strong model. The standard test set which contains 1k data samples was used for the test. The data is restructured into a balanced binary classification task, i.e., given a question and an answer, the model is required to determine whether the answer is correct.
@@
Slot filling. The performance of WeakS-to-Strong on the reliability of generated content was evaluated on the slot-filling task, which is a crucial spoken language understanding task aiming at filling in the correct value for predefined slots (e.g. restaurant and hotel names). SLURP dataset (Bastianelli et al., 2020) was used which contains 16.5k sentences and 72k audio recordings of single-turn user interactions with a home assistant, annotated with scenarios, actions and entities. Only the reference transcriptions of the speech were used for training. Following Sun et al. (2023a; b), we designed the prompt with slot keys and descriptions in the same way. In our setup, 2k utterances from the train split were extracted for training the weak models, and another 2k utterances were reserved for generating weak labels and training the strong model. We report the performance of both weak and strong models on the standard SLURP test set.
5.2 Models and baselines
Models. For classification, the Qwen-7B (Bai et al., 2023)222https://huggingface.co/Qwen/Qwen-7B model was applied as the strong student model. Five models were used as weak teachers: GPT2-Large (Radford et al., 2019)333https://huggingface.co/openai-community/gpt2-large, OPT-1.3B (Zhang et al., 2022)444https://huggingface.co/facebook/opt-1.3b, Pythia-1.4B (Biderman et al., 2023)555https://huggingface.co/EleutherAI/pythia-1.4b, BLOOM-1B1 (Le Scao et al., 2022)666https://huggingface.co/bigscience/bloom-1b1, and TinyLlama_v1.1 (Zhang et al., 2024)777https://huggingface.co/TinyLlama/TinyLlama_v1.1. The last linear layer which maps the embeddings to tokens is replaced with a linear classification head with two outputs to adapt language models to the classification setting.